OpenStudy (anonymous):

Solve the equation 2x - 7 = 5

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

2x - 7 +7 = 5 +7 2x = 12 x = 12/2 x = 6

OpenStudy (anonymous):

Wooooooops!!!

OpenStudy (anonymous):

2x - 7 +7 = 5 +7 2x = 12 x = 12/2 = 6

OpenStudy (anonymous):

Add 7 to both sides this will leave you with 2x=12. then divide both sides by 2 leaving you with an answer of x=6